OpenStudy (anonymous):

Why can you use cross products to solve the proportion 18/5 = x/100 for x? A because 18 times 100 is equal to 5 times x, which means that x is equal to 36 B because 18 times 100 is equal to 5 times x, which means that x is equal to 360 C because 18 times x is equal to 5 times 100, which means that x is equal to 36 D because 18 times x is equal to 5 times 100, which means that x is equal to 360

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@misty1212

OpenStudy (cathyangs):

Do you know what "cross products" mean? It is what it sounds like, finding the product using the cross. So for A/B = C/D, using cross products you get A * D = B *C You can think of the lines of multiplication making a cross over the equals sign in the proportion. Does this make sense?

OpenStudy (anonymous):

i know cross products i didn't even think about them thx for reminding me

OpenStudy (cathyangs):

No problem! What do you think the answer is, then? You basically just need to solve the proportion for x using cross products, then see which statement matches your work.

OpenStudy (anonymous):

i think it's D

OpenStudy (cathyangs):

Not quite. Think about the equation I said before: for A/B = C/D, using cross products you get A * D = B *C What would the A * D = B * C equation be, using the values in your problem?

OpenStudy (anonymous):

it would be 18 x 100 and 5 x X

OpenStudy (cathyangs):

Correct!

OpenStudy (cathyangs):

That fits with choices A and B. How can you simplify that to find what x is?

OpenStudy (anonymous):

so its b

OpenStudy (cathyangs):

Correct! Good work.

OpenStudy (anonymous):

thx heres the last one

OpenStudy (anonymous):

never mind i don't need help on the last one i'll follow last questions steps thank you very much for helping me

OpenStudy (cathyangs):

No problem! Here to help :)